OpenStudy (anonymous):

Solve for \(x\) \[2L_2 x^3 =2L_{1}(m-x)^3\]

Parth (parthkohli):

What do \(\rm L_1\) and \(\rm L_2\) stand for?

Parth (parthkohli):

Dividing both sides by 2,\[\rm L_2 x^3 =L_1(m - x)^3\]

OpenStudy (anonymous):

You don't need to know what they are..this is a minor problem from a larger problem..you can just treat it as some random variable.. I just need \(x\)

Parth (parthkohli):

Okey dokey.

OpenStudy (anonymous):

And you need to remind me how to solve it..I don't seem to get it; to continue solving this problem

Parth (parthkohli):

\[\rm L_2 x^3=L_1(m^3 -3m^2x+3mx^2 - x^3)\]I'm just gonna say that L1 = m and L2 = n

Parth (parthkohli):

no. L1 = a and L2 = b.

Parth (parthkohli):

\[\rm bx^3 = a(m^3 - 3m^2x + 3mx^2 - x^3)\]

OpenStudy (anonymous):

Oh no. I want a simple way please!

OpenStudy (anonymous):

It's some algebra law/ rulles that i forgot

Parth (parthkohli):

Simple way? http://wolframalpha.com Enjoy!

OpenStudy (anonymous):

Lol, doesn't work. That is why I'm here :P

Parth (parthkohli):

Step by step solution. \(\rm \checkmark \)

OpenStudy (anonymous):

lol, does not work..it doesn't compute

Parth (parthkohli):

The way I was doing it is actually the simplest. lol

OpenStudy (anonymous):

Not really :)\[L_2 x^3 =L_{1}(m-x)^3 =>\sqrt[3]{L_{2}}x=\sqrt[3]{L_{1}} (m-x)\] But i dont get what rules it comes froms?

Parth (parthkohli):

Oh, it's just finding the cube root of both sides.

Parth (parthkohli):

HOW DIDN"T THAT COME TO MY MIND? lol

OpenStudy (anonymous):

lol, i cant remmember how to do this

OpenStudy (anonymous):

so just assume that you can cube root both sides?

Parth (parthkohli):

@Omniscience Are you seriously serious? You don't know how to cube root both sides

Parth (parthkohli):

You can always cube root both sides. -.-

Parth (parthkohli):

Whatever you do, do the same to both sides.

OpenStudy (anonymous):

lol..i forgot..but i thought it was only one side?

Parth (parthkohli):

Both sides. BOTH SIDES ALWAYS

Parth (parthkohli):

You gotta maintain the balance.

OpenStudy (anonymous):

\[L_2 x^3 =L_{1}(m-x)^3 => L_2 = \sqrt[3]{L_1(m-x)^2}\] i thought you have to do it like this..

Parth (parthkohli):

But also remember,\[\sqrt[3]{ab} = \sqrt[3]{a} \times \sqrt[3]{b}\]

OpenStudy (anonymous):

I know that..but why do you have to cube roots both sides? it just not make sense lol

Parth (parthkohli):

BECAUSE BOTH SIDES ARE EQUAL.

OpenStudy (anonymous):

lol..can you do the for sqrt as well?

Parth (parthkohli):

Yeah, all radicals.

OpenStudy (anonymous):

srsly!!!

Parth (parthkohli):

Yeah. -.-

OpenStudy (anonymous):

I feel like an idiot lol

OpenStudy (anonymous):

:'(